OpenStudy (anonymous):

PLEASE HELP!!! Which of the following is a polynomial with roots -2, -3i, and 3i ? a.) x3 + 4x2 + 9x + 24 b.) x3 - 4x2 + 9x - 24 c.) x3 + 2x2 + 9x + 18 d.) x3 - 2x2 + 9x - 18

OpenStudy (jdoe0001):

so, you have the roots x = -2 x= -3i x = 3i so that means, x+2 = 0 x+3i = 0 x-3i = 0 so (x+2)(x+3i)(x-3i) =0 there's your polynomial, expand the roots

OpenStudy (anonymous):

another way to do it is from x = -2 x = -3i x = 3i you simplify by squaring those complex roots x = -2 x^2 = -3(-1) x^2 = 3(-1) so the roots would be x = -2 x^2 = 3 x^2 = -3 so the factored form is (x + 2)(x^2 - 3)(x^2 + 3) whichever you think looks easier

OpenStudy (anonymous):

i got c as the answer